### v1

@@ Abstract
A central challenge in neuroscience is understanding how the brain’s complex nonlinear dynamics give rise to emergent properties underlying adaptive cognition and behavior across spatial and temporal scales. These dynamics exhibit scale-free and multifractal properties, shaping the adaptive reconfiguration of neural networks. However, conventional neuroimaging models remain constrained by assumptions of linearity and stationarity, limiting their capacity to fully characterize these intricate processes. Transformer-based architectures, known for capturing long-range dependencies in sequential data, naturally align with the brain’s intrinsically hierarchical and temporal organization. Here, we introduce Multi-Band Brain Net (MBBN), a transformer-based deep learning framework that explicitly models frequency-specific spatiotemporal brain dynamics from fMRI by integrating scale-free network principles with frequency-resolved multi-band self-attention mechanisms.
Trained on three large-scale neuroimaging cohorts (UK Biobank, ABCD, ABIDE) totaling 45,951 individuals, MBBN reveals previously undetectable frequency-dependent intrinsic network interactions, shedding new light on distinct connectivity disruptions across psychiatric conditions (ADHD, ASD, depression). This extensive validation demonstrates robust generalizability and highlights core neural principles conserved across populations and developmental stages.
Specifically, MBBN achieves up to 30.59% higher predictive accuracy than state-of-the-art deep learning and conventional connectome-based methods, demonstrating the advantage of frequency-informed spatiotemporal modeling in capturing latent neural computations. Furthermore, MBBN’s attention-based interpretability reveals novel frequency-specific biomarkers for neurodevelopmental disorders, offering mechanistic insights into the hierarchical organization of brain function.
By providing a principled, interpretable framework for dissecting spatiotemporal learning mechanisms, MBBN advances methodological rigor and provides new insight into how dynamic neural computations underpin cognitive function and psychiatric vulnerability, with implications for brain decoding, cognitive neuroscience, and precision psychiatry.
@@ 1 Introduction
Understanding the brain’s dynamic processes is essential for unraveling the mechanisms underlying adaptive cognition and behavior. These processes influence critical functions such as learning, decision-making, and social interactions, and their dysregulation is closely linked to mental health conditions ([1, 2, 3, 4]). While functional magnetic resonance imaging (fMRI) has been instrumental in probing brain activity, many conventional analytical methods remain constrained by assumptions of linearity, stationary, static connectivity. Approaches ranging from mass-univariate statistical tests and seed-based correlations to static functional connectivity maps often treat the brain as a relatively fixed system, thus failing to capture the temporal fluctuations, nonlinear interactions, and adaptive reconfigurations that characterize real neural dynamics. Recent advances in artificial intelligence (AI), particularly transformer-based architectures, have begun to address these shortcomings by modeling brain activity as a continuously evolving network of interdependent processes ([5, 6, 7, 8]).
However, a critical gap remains in understanding and modeling two fundamental properties of neural dynamics: scale-free dynamics and multifractal structures. These properties are essential for comprehending the brain’s hierarchical complexity and adaptability ([9, 10, 2]). Scale-free dynamics, characterized by power-law relationships (Equation 1), reveal how neural activity maintains consistent patterns across different timescales and modalities. These dynamics play crucial roles in brain maturation and state transitions, with temporal variability serving as a key regulatory mechanism ([11, 12]). The power-law exponent ($\beta$) varies across brain regions and states, providing valuable insights into neural network heterogeneity and serving as a potential marker for neurological and psychiatric disorders ([13, 14]).
Multifractality extends this framework by incorporating multiple scaling dimensions and power-law behaviors ([15, 16, 17]), offering a sophisticated approach to understanding the brain’s nested, hierarchical organization ([18, 19]). This property manifests in distinct frequency-dependent patterns that reflect neural adaptability and functional specialization ([20]). Evidence from fMRI, ECoG, and EEG studies suggests that these multifractal principles drive frequency-specific changes in connectivity and topological organization, particularly in properties such as small-worldness and modularity ([21, 22]). Understanding these patterns is crucial for developing more effective diagnostic and therapeutic strategies.
@@
To address these challenges, we introduce the Multi-Band Brain Net (MBBN), a novel framework based on Bidirectional Encoder Representations from Transformers (BERT) that integrates scale-free property-based frequency decomposition with advanced self-attention mechanisms and network theory concepts like communicability (Figure 1). MBBN divides fMRI BOLD signals into distinct frequency bands, enabling the capture of band-specific functional and topological properties. This approach provides a neurobiologically informed framework for modeling the brain’s multifractal and scale-free dynamics. The model features two key strengths: Firstly, its self-attention weighted connectivity dynamically models functional connectivity by capturing complex, nonlinear interactions and time-varying relationships between brain regions, thereby overcoming the limitations of traditional static, correlation-based linear representations. Secondly, a communicability-inspired pretraining loss function optimizes for biologically meaningful connectivity patterns. By leveraging these advances, MBBN not only provides deeper insights into brain function but also identifies frequency-specific markers for neurological and psychiatric disorders, offering a powerful new tool for clinical neuroscience research and applications.
@@
The MBBN module processes each timeseries signal, divided by frequency, within specific frequency ranges. In the temporal module (denoted as BERT), BERT updates the [CLS (Classification)] token, which is then passed through a classifier to generate a prediction $\hat{y_{\nu}}$ ($\nu$ represents the frequency range). The spatial module employs a multi-head attention mechanism to compute attention matrices representing the interactions between regions of interest (ROIs). B. The MBBN modules are applied to timeseries signals across three frequency ranges (ultra-low, low, and high frequency). Parameter sharing is implemented in the temporal module, while spatial modules have independent parameters for each frequency range. The final output of MBBN is computed as the average of the predictions across all frequency ranges, along with attention matrices specific to each frequency range. C. The frequency decomposition process is demonstrated using a representative sample from the UKB dataset. Although the knee frequencies ($f_{1}$, $f_{2}$) may vary across individuals, the overall pattern of frequency division remains consistent. The blue line represents the original power spectral density (PSD) of the fMRI signal. The pink, yellow-green, and yellow lines show the fitted power-law functions for the high, ultra-low, and low-frequency ranges, respectively. The dotted red and green lines indicate the frequency boundaries: ultra-low to low ($f_{1}$), and low to high ($f_{2}$) frequencies, respectively.
@@ 2 Results
2.1 Comparison of Model Performance Across Various Downstream Tasks
@@
Figure 2: Comparative Analysis of fMRI Prediction Models. Across datasets and tasks, the proposed MBBN demonstrates superior predictive performance compared to baseline models including state-of-the-art models in the field (e.g., Brain Network Transformer ([5])), as well as vanilla BERT (without multi-fractality division). [CC] indicates models based on Pearson’s correlation-based connectivity; [TD] represents, models accounting for temporal dynamics. The x-axis denotes the models, and the y-axis represents performance metrics, with AUROC used for classification tasks and MAE for regression tasks. (A–B) show results for sex prediction, (C–D) for fluid intelligence prediction, (E–F) for depression prediction, G for ADHD prediction, and H for ASD prediction. Specifically: (A, B, C, G) correspond to the ABCD dataset using the HCP-MMP1 atlas. (D, E, F) correspond to the UKB dataset using the HCP-MMP1 atlas. H correspond to the ABIDE dataset using the HCP-MMP1 atlas.
To evaluate model performance, we trained MBBN from scratch on multiple downstream tasks spanning biological, cognitive, and clinical domains, using both the HCP-MMP1 and Schaefer atlases across the UKB and ABCD datasets. MBBN consistently outperformed baseline models across all tasks. For biological phenotype prediction (sex classification), MBBN achieved notable gains of up to 11.6% on UKB and 18.88% on ABCD using the HCP-MMP1 atlas, with similar improvements using the Schaefer atlas (9.06% on UKB, 14.78% on ABCD). In cognitive assessment tasks, fluid intelligence prediction showed substantial improvements on the ABCD dataset (up to 18.88% with HCP-MMP1, 8.03% with Schaefer). For clinical applications, depression prediction on the UKB dataset demonstrated meaningful gains (up to 9.42% with HCP-MMP1, 6.98% with Schaefer). Comprehensive performance comparisons, including task-specific metrics and statistical analyses, are detailed in Appendix.
2.2 Communicability-based masking loss improves fine-tuning performance over random masking loss.
@@
A key component of our methodology was a pretraining loss function designed to mask nodes with high communicability within the brain network. This strategy was predicated on the hypothesis that selectively reconstructing these highly influential nodes during pretraining would better represent the complex communication patterns within the brain. To validate this hypothesis and evaluate the downstream impact of this pretraining strategy, MBBN was subsequently fine-tuned on two clinically relevant classification tasks: ADHD classification using the ABCD dataset, and ASD classification using the ABIDE dataset. These datasets and tasks were chosen for their established clinical significance and high prevalence rates.
@@
Fine-tuning experiments revealed a clear advantage for the MBBN-high model, which employed a pretraining strategy that selectively masked nodes with high communicability. This approach resulted in significant performance improvements compared to baselines: a 5.55% gain in ADHD classification accuracy and a remarkable 30.59% gain in ASD classification accuracy. In contrast, models using random node masking (MBBN-random) or masking nodes with low communicability (MBBN-low) showed inferior performance, even compared to the MBBN model trained from scratch (MBBN-from scratch). These findings underscore the critical role of communicability-based pretraining in effectively leveraging the underlying network structure of the data to enhance predictive performance. Complete results and statistical details are available in Appendix.
2.3 Key connectivity for mental illness classification revealed by attention mechanisms
Figure 3: Self-Attention-Based Connectivity Learned from Resting FMRI Distinguishing ADHD and ASD from Healthy Controls. For presentational purposes, this figure highlights statistically significant connections with the top 10 largest effect sizes that differentiate individuals with ADHD (A-C) or ASD (D-F) from healthy controls (HC) (FDR corrected P’s ¡ 0.05). [Red: Higher in the disorder group; Blue: Lower in the disorder group.]. A detailed list of the top 100 connectivity is provided in the Appendix F. (A-C) show ADHD connectivity patterns compared to HC across frequency bands: high (A), low (B), and ultralow (C). (D-F) illustrate ASD patterns compared to HC: high (D), low (E), and ultralow (F). The HCP-MMP1 atlas was used.
We utilized a Grad-CAM-based attribution score to identify self-attention-weighted-connectivity with large effect sizes distinguishing healthy controls from individuals with mental illness and analyzed their roles. Details are described in Section 3.9.
@@ 2.3 Key connectivity for mental illness classification revealed by attention mechanisms
In the ABCD test dataset (HCP-MMP1 atlas), frequency-dependent differences in functional connectivity patterns were identified between the ADHD and HC groups. ADHD showed reduced connectivity in high and low frequencies, while exhibiting increased connectivity in the ultralow-frequency range (Figure 3). These findings reveal frequency-specific neural connectivity alterations in ADHD. The results of the t-test distinguishing between the HC and ADHD groups showed that 6.55% of the statistically significant connections ($p<0.05$) were in the high-frequency range, with the 75% quantile value of Cohen’s d score being 0.71. In contrast, the low-frequency range accounted for only 1.67%, with the 75% quantile value of Cohen’s d being 0.17. The ultralow frequency range accounted for 8.39%, with the 75% quantile value of Cohen’s d being 0.71.
In the high-frequency range (Figure 3 A), distinct connectivity patterns emerged involving key regions such as ParaHippocampal Area 3 and VentroMedial Visual Area, highlighting potential disruptions in visual processing and memory formation. Similarly, the low-frequency range (Figure 3 B) showed significant differences centered around Area PH (Parahippocampal Area) and Primary Auditory Cortex , pointing to altered neural system how sound information is integrated and processed in memoryn. In the ultralow-frequency range (Figure 3 C), connectivity changes in VentroMedial Visual Area 3 and Area TF suggest impairments related to visual processing and attention regulation.
In the ABIDE test dataset (HCP-MMP1 atlas), the results of the t-test distinguishing between the HC and ASD groups showed that 7.88% of the statistically significant connections ($p<0.05$) were in the high-frequency range, with the 75% quantile value of Cohen’s d score being 1.19. In contrast, the low-frequency range accounted for only 0.24%, with the 75% quantile value of Cohen’s d being 0.74. The ultralow frequency range accounted for 0.91%, with the 75% quantile value of Cohen’s d being 0.84.
ASD-related connectivity patterns exhibited localized differences compared to healthy controls (HC), with varying effects across frequency ranges (Figure 3). In the high-frequency range (Figure 3 D), connectivity differences were observed in regions such as PGi (PG inferior) and 47m (Area 47 medial), potentially linked to challenges in socio-emotional and executive functions. In the low-frequency range (Figure 3 E), IFSa showed prominent differences, indicating possible impairments in executive and communication functions. Finally, in the ultralow-frequency range (Figure 3 F), connectivity alterations in Lateral Area 7P and the Insular Granular Complex suggest disruptions in sensory and emotional processing.
In both the ABCD and ABIDE datasets, analysis of false positive and false negative subjects (based on predicted values) showed no statistically significant connections in the self-attention-based connectivity analysis.

### v2

@@ Abstract
Understanding how the brain’s complex nonlinear dynamics give rise to cognitive function remains a central challenge in neuroscience. While brain functional dynamics exhibits scale-free and multifractal properties across temporal scales, conventional neuroimaging analytics assume linearity and stationarity, failing to capture frequency-specific neural computations. Here, we introduce Multi-Band Brain Net (MBBN), the first transformer-based framework to explicitly model frequency-specific spatiotemporal brain dynamics from fMRI. MBBN integrates biologically-grounded frequency decomposition with multi-band self-attention mechanisms, enabling discovery of previously undetectable frequency-dependent network interactions. Trained on 49,673 individuals across three large-scale cohorts (UK Biobank, ABCD, ABIDE), MBBN sets a new state-of-the-art in predicting psychiatric and cognitive outcomes (depression, ADHD, ASD), showing particular strength in classification tasks with up to 52.5% higher AUROC and provides a novel framework for predicting cognitive intelligence scores. Frequency-resolved analyses uncover disorder-specific signatures: in ADHD, high-frequency fronto-sensorimotor connectivity is attenuated and opercular somatosensory nodes emerge as dynamic hubs; in ASD, orbitofrontal-somatosensory circuits show focal high-frequency disruption together with enhanced ultra-low-frequency coupling between the temporo-parietal junction and prefrontal cortex. By integrating scale-aware neural dynamics with deep learning, MBBN delivers more accurate and interpretable biomarkers, opening avenues for precision psychiatry and developmental neuroscience.
@@ 1 Introduction
Understanding the brain’s dynamic processes is essential for unraveling the mechanisms underlying adaptive cognition and behavior. These processes influence critical functions such as learning, decision-making, and social interactions, with mounting evidence that their dysregulation is closely linked to mental health conditions ([1, 2, 3, 4]). While functional magnetic resonance imaging (fMRI) has been instrumental in probing brain activity, many conventional analytical methods remain constrained by assumptions of linearity, stationarity and static connectivity. Approaches ranging from mass-univariate statistical tests and seed-based correlations to static functional connectivity maps often treat the brain as a relatively fixed system, thus failing to capture the temporal fluctuations, nonlinear interactions, and adaptive reconfigurations that characterize real neural dynamics. Recent advances in artificial intelligence, particularly transformer-based architectures originally designed for sequential data, have shown promise for modeling brain activity as a continuously evolving network of interdependent processes. Yet, these approaches largely ignore the frequency-specific nature of neural oscillations ([5, 6, 7, 8]).
This oversight is particularly problematic given that existing neuroimaging models, including recent deep learning approaches, treat brain signals as broadband phenomena, overlooking the fundamental frequency-specific organization of neural computation. Indeed, this organization is best understood through two fundamental properties that describe the brain’s complexity: its multifractal structure and its scale-free dynamics.
Multifractality, a key structural property, points to the brain’s nested, hierarchical organization where multiple scaling dimensions and power-law behaviors coexist ([9, 10, 11, 12, 13]). Crucially, it is within this complex structure that scale-free dynamics—a key functional property—become apparent. While a broadband signal is not inherently scale-free, decomposing it into distinct frequency components reveals power-law relationships (Equation 1) that signify consistent activity patterns across different timescales. These dynamics play crucial roles in brain maturation and state transitions, with temporal variability serving as a key regulatory mechanism ([14, 15]). The power-law exponent ($\beta$) varies across brain regions and states, providing valuable insights into neural network heterogeneity and serving as a potential marker for neurological and psychiatric disorders ([16, 17]). Evidence from fMRI, ECoG, and EEG studies further suggests that these principles drive frequency-specific changes in connectivity and topological organization, such as small-worldness and modularity ([18, 19]). Understanding these patterns is crucial for developing more effective diagnostic and therapeutic strategies.
@@
To address these challenges, we introduce Multi-Band Brain Net (MBBN), the first transformer-based framework specifically designed to model frequency-resolved spatiotemporal brain dynamics from fMRI. Unlike existing approaches that analyze broadband signals, MBBN explicitly decomposes neural activity into biologically meaningful frequency bands using scale-free principles, then applies specialized attention mechanisms to capture frequency-specific connectivity patterns incorporating network theory concepts like communicability (Figure 1, 4). MBBN divides fMRI BOLD signals into distinct frequency bands, enabling the capture of band-specific functional and topological properties. This approach provides a neurobiologically informed framework for modeling the brain’s multifractal and scale-free dynamics. The model features two key strengths: Firstly, its self-attention weighted connectivity dynamically models functional connectivity by capturing complex, nonlinear interactions and time-varying relationships between brain regions, thereby overcoming the limitations of traditional static, correlation-based linear representations. Secondly, a communicability-inspired pretraining loss function optimizes for biologically meaningful connectivity patterns. By leveraging these advances, MBBN not only provides deeper insights into brain function but also identifies frequency-specific markers for neurological and psychiatric disorders, offering a powerful new tool for clinical neuroscience research and applications.
@@
The MBBN module processes each timeseries signal, divided by frequency, within specific frequency ranges. In the temporal module (denoted as BERT), BERT updates the [CLS (Classification)] token, which is then passed through a classifier to generate a prediction $\hat{y_{\nu}}$ ($\nu$ represents the frequency range). The spatial module employs a multi-head attention mechanism to compute attention matrices representing the interactions between regions of interest (ROIs). B. The MBBN modules are applied to timeseries signals across three frequency ranges (ultra-low, low, and high frequency). Parameter sharing is implemented in the temporal module, while spatial modules have independent parameters for each frequency range. The final output of MBBN is computed as the average of the predictions across all frequency ranges, along with attention matrices specific to each frequency range. C. Frequency decomposition using scale-free principles on representative UKB data. Individual-specific knee frequencies ($f_{1}$, $f_{2}$) separate three biologically distinct bands: ultralow (pink), low (yellow-green), and high (yellow) frequencies, each exhibiting characteristic power-law scaling. The dotted red and green lines indicate the frequency boundaries: ultra-low to low ($f_{1}$), and low to high ($f_{2}$) frequencies, respectively.
@@ 2 Results
2.1 MBBN demonstrates superior performance across diverse neuroimaging tasks
@@
Figure 2: Comparative Analysis of fMRI Prediction Models. Across datasets and tasks, the proposed MBBN demonstrates superior predictive performance compared to baseline models including state-of-the-art models in the field (e.g., Brain Network Transformer ([5])), as well as vanilla BERT (without multifractal decomposition). [CC] indicates models based on Pearson’s correlation-based connectivity; [TD] represents models accounting for temporal dynamics. The x-axis denotes the models, and the y-axis represents performance metrics, with AUROC used for classification tasks and MAE for regression tasks. MBBN consistently outperforms baseline models across biological (A-B: sex prediction), cognitive (C-D: fluid intelligence), and clinical (E-H: depression, ADHD, ASD) prediction tasks. [CC] indicates correlation-based connectivity models; [TD] represents temporal dynamics models. AUROC is shown for classification tasks, MAE for regression tasks. Data from ABCD (A-C, G), UKB (D-F), and ABIDE (H) datasets using HCP-MMP1 atlas. A-F denotes from-scratch experiments and G-H denotes fine-tuning experiments.
Task MBBN vs Best Baseline (%) vs Worst Baseline (%)
ABCD Sex 0.916 (AUROC) +2.346 +19.271
ABCD Fluid Intelligence 0.717 (MAE) +0.693 +9.584
ABCD Depression 0.556 (AUROC) +1.460 +7.961
UKB Sex 0.980 (AUROC) +3.158 +14.086
UKB Fluid Intelligence 1.630 (MAE) -3.099 -1.305
UKB Depression 2.21 (MAE) -1.231 +0.924
ABCD ADHD 0.645 (AUROC) +1.735 +7.860
ABIDE ASD 0.810 (AUROC) +24.615 +52.542
Table 1: Performance gain of MBBN between best and worst baseline model. We illustrated finetuned MBBN for comparison of ABCD ADHD classification and ABIDE ASD classification.
To evaluate its performance, MBBN was trained from scratch on a diverse set of downstream tasks spanning biological, cognitive, and clinical domains, utilizing both the HCP-MMP1 and Schaefer atlases with the UKB and ABCD datasets. While MBBN demonstrated unparalleled performance in classification tasks, its advantage in regression tasks (e.g., fluid intelligence and depression score prediction) was less pronounced, suggesting that the model’s architecture is particularly optimized for distinguishing categorical phenotypes rather than predicting continuous scores. The primary results presented in Figure 2 A-F focus on experiments conducted with the HCP-MMP1 atlas; corresponding analyses using the Schaefer atlas are detailed in the Appendix E.1.
For biological phenotype prediction, MBBN achieved substantial improvements in sex classification with gains up to 14.09% (UKB) and 19.27% (ABCD) using HCP-MMP1 atlas. Cognitive assessment showed marked improvements, with fluid intelligence prediction demonstrating up to 9.58% enhancement on ABCD. Also clinical prediction tasks revealed MBBN’s therapeutic potential, achieving 0.92% improvement in depression prediction (UKB) and exceptional performance in psychiatric disorders (detailed in Supplementary Materials).
2.2 Communicability-based pretraining enables superior clinical prediction performance
@@
The exceptional performance of MBBN, particularly in clinical prediction tasks, stems from our novel communicability-based pretraining strategy that masks nodes with high communicability within the brain network. This strategy was predicated on the hypothesis that selectively reconstructing these highly influential nodes during pretraining would better represent the complex communication patterns within the brain. To validate this hypothesis and evaluate the downstream impact of this pretraining strategy, MBBN was subsequently fine-tuned on two clinically relevant classification tasks: ADHD classification using the ABCD dataset, and ASD classification using the ABIDE dataset. These datasets and tasks were chosen for their established clinical significance and high prevalence rates.
@@
Fine-tuning experiments (Figure 2 G-H) revealed a clear advantage for the MBBN-high model, which employed a pretraining strategy that selectively masked nodes with high communicability. This approach resulted in significant performance improvements compared to baselines: a 7.86% gain in ADHD classification AUROC and a 52.54% gain in ASD classification AUROC. In contrast, models using random node masking (MBBN-random) or masking nodes with low communicability (MBBN-low) showed inferior performance, even compared to the MBBN model trained from scratch (MBBN-from scratch). These findings underscore the critical role of communicability-based pretraining in effectively leveraging the underlying network structure of the data to enhance predictive performance. Detailed results and statistical details are available in Appendix E.1. These results demonstrate that communicability-based loss better captures a complex information flow of the brain than conventional masking loss, and MBBN’s frequency-resolved approach captures clinically relevant neural signatures that are missed by conventional broadband analysis methods.
2.3 Frequency-specific connectivity patterns reveal distinct psychiatric biomarkers
Figure 3: Frequency-specific connectivity disruptions in ADHD and ASD. Statistically significant connections (FDR-corrected $p<0.05$, top 10 effect sizes) distinguishing ADHD (A-C) and ASD (D-F) from healthy controls across high, low, and ultralow frequency bands. Red: increased connectivity in disorder; Blue: decreased connectivity. Detailed results in Supplementary Materials.
To identify frequency-specific biomarkers underlying psychiatric disorders, we applied Grad-CAM-based attribution analysis to extract self-attention-weighted connectivity patterns that distinguished patient groups from healthy controls. This approach revealed distinct frequency-dependent neural signatures for ADHD and ASD that were previously undetectable using conventional broadband connectivity analysis. Details are described in Section 3.9.
@@ 2.3 Frequency-specific connectivity patterns reveal distinct psychiatric biomarkers
In the ABCD test dataset (HCP-MMP1 atlas), frequency-dependent differences in functional connectivity patterns were identified between the ADHD and HC groups. ADHD showed reduced connectivity in high and low frequencies, while exhibiting increased connectivity in the ultralow-frequency range (Figure 3). These findings reveal frequency-specific neural connectivity alterations in ADHD.
ADHD exhibits frequency-dependent connectivity alterations with high-frequency disruptions predominating. Comparing ADHD patients to healthy controls in the ABCD dataset, we identified distinct patterns of connectivity changes across frequency bands (Figure 3 A-C). Statistical analysis revealed that high-frequency connections showed the most extensive alterations, with 8.43% of all significant connections ($p<0.05$) occurring in this band and a 75th percentile Cohen’s d of 0.332, indicating moderate effect sizes. In contrast, low-frequency and ultralow-frequency bands showed more limited but still significant changes, accounting for 3.46% (Cohen’s d 75th percentile = 0.220) and 3.56% (Cohen’s d 75th percentile = 0.117) of significant connections, respectively.
Anatomically, high-frequency disruptions centered on Area 10v (ventral part of Brodmann Area 10), indicating impaired executive control networks. Low-frequency alterations consistently involved Area OP2-3-VS (opercular parts of Areas 2 and 3 and adjacent Ventral Somatosensory / Sylvian region) and Auditory 4 Complex, suggesting disrupted sensorimotor-auditory integration critical for attention regulation. Ultralow-frequency changes in Medial Belt Complex and ParaHippocampal regions, while showing smaller effect sizes, point to subtle alterations in memory and emotional processing circuits.
ASD demonstrated a contrasting frequency profile with ultralow and low frequencies most severely affected (Figure 3 D-F). T-test comparisons between ASD patients and healthy controls in the ABIDE dataset revealed a different pattern from ADHD. Ultralow-frequency connections showed the most extensive disruptions, representing 4.54% of all significant connections with large effect sizes (75th percentile Cohen’s d = 0.717). Low-frequency alterations were similarly prominent, accounting for 4.04% of significant connections with substantial effect sizes (Cohen’s d 75th percentile = 0.714). Notably, high-frequency changes were minimal, comprising only 1.81% of significant connections, though with moderate effect sizes (Cohen’s d 75th percentile = 0.575).
High-frequency changes, while limited in extent, localized specifically to Area 11l (lateral part of Brodmann Area 11), consistent with known social-cognitive deficits. Low-frequency disruptions prominently affected Area 10pp (frontopolar cortex), suggesting impaired higher-order executive functions. Most significantly, the extensive ultralow-frequency alterations involved Area TPOJ3 (temporo-parieto-occipital junction) and multiple frontal regions, indicating fundamental disruptions in social cognition and language networks—core features of ASD pathophysiology.
In both the ABCD and ABIDE datasets, analysis of false positive and false negative subjects (based on predicted values) showed no statistically significant connections in the self-attention-based connectivity analysis. These frequency-resolved patterns reveal disorder-specific neural signatures: ADHD primarily affects high-frequency sensorimotor networks with compensatory ultralow-frequency changes, while ASD shows pervasive low and ultralow-frequency disruptions in social-cognitive circuits. Importantly, false positive and false negative cases showed no significant connectivity patterns, confirming the specificity of these frequency-dependent biomarkers for accurate psychiatric classification.